# Congruences for the ratios of Rankin–Selberg $L$-functions

### v1

@@ 2.3.2. The algorithm
    For a given eigenform $f\in S_{k}(\Gamma_{1}(N))$ construct the space of modular symbols $\mathbb{M}_{k}(\Gamma_{1}(N)).$
@@
    Find a Manin basis for $\mathbb{M}_{k}(\Gamma_{1}(N))$. Call them $e_{1},e_{2},\dots,e_{n}$. Express the modular symbols $b_{i}:=X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\}$ for $i=1,2,\dots,k-1$ in terms of the basis $e_{1},e_{2},\dots,e_{n}$. Say $b_{i}=\sum_{j=1}^{n}a_{ji}e_{j}$ for $i=1,2,\dots,k-1$. Let $A$ be the matrix $a_{ij}]$.
@@
    For a prime $p$ find the matrix of the $p$-th Hecke operator $T_{p}$ with respect to the basis $e_{1},e_{2},\cdots e_{n}$. Call that matrix to be $B=[b_{ij}].$
@@
    Form a matrix $\mathcal{L}$ as follows:
    $\mathcal{L}=\left(\begin{array}[]{c|c}I_{k-1}&A^{t}\\ \hline\cr\\ \mathbf{0}_{n,k-1}&B-a(p,f)\cdot I_{n}\end{array}\right).$
@@
    Perform row reduction on $\mathcal{L}$. Call the resulting matrix $\mathcal{L}_{\text{red}}$. For the values $i=1,3,5\cdots$, find a column in $\mathcal{L}_{\text{red}}$ such that the first entry of the column is non-zero. Let such a column be $C_{\text{odd}}=(c_{1},c_{2},\cdots,c_{n+k-1})^{t}$. So $c_{1}\neq 0.$ Similarly, find a column $C_{\text{even}}=(d_{1},d_{2},\cdots,d_{n+k-1})^{t}$ of $\mathcal{L}_{\text{red}}$ such that $d_{2}\neq 0$.
@@
    Return: $\dfrac{1}{(i-1)!}\cdot\dfrac{c_{i}}{c_{1}}\text{ for $i=1,3,5,\cdots$ and }\dfrac{1}{(i-1)!}\cdot\dfrac{d_{i}}{d_{2}}\text{ for $i=2,4,\cdots$}.$
This is the desired value due to
$\frac{(f,X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\})}{(f,Y^{k-2}\otimes\{0,\infty\})}=\frac{c_{i}}{c_{1}}\>\>\>\mbox{for}\>\>i=1,3,5\cdots,$
$\frac{(f,X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\})}{(f,XY^{k-3}\otimes\{0,\infty\})}=\frac{d_{i}}{d_{2}}\>\>\>\mbox{for}\>\>i=2,4\cdots$
and the equations (11) and (12).
2.3.3. An example
Suppose $f$ is a newform of weight $5$ for $\Gamma_{1}(4)$ with $q$-expansion is $f(q)=q-4q^{2}+16q^{4}+\cdots$. The algorithm in 2.3.2 when used to compute the special values of $D(s,f)$ gives us:
$\frac{D(4,f)}{D(2,f)}=\frac{\pi^{2}}{5}\ \ \text{ and }\ \ \frac{D(3,f)}{D(1,f)}=\frac{5\pi^{2}}{12}.$
The interested reader is referred to the first author’s thesis [10] for details of the implementation of the algorithm.
@@ 2.4. Sturm’s criterion for congruence between cusp forms
To check congruence between cusp forms we will use the following criterion of Sturm [18]; see also Stein [17, p. 173].
@@
Let $\mathfrak{P}$ be a prime ideal of a number field $K$. Let $f,g\in M_{k}(\Gamma,\mathcal{O}_{K})$ be modular forms of weight $k$ for a congruence subgroup $\Gamma$ of level $N$ with Fourier coefficients in the ring of integers $\mathcal{O}_{K}$ of $K$. Let $m=[SL_{2}(\mathbb{Z}):\Gamma].$ Suppose
@@
$n\leq\begin{cases}\frac{km}{12}-\frac{m-1}{N}&\mbox{if $f-g\in S_{k}(\Gamma,\mathcal{O}$}),\\ \frac{km}{12}&\mbox{otherwise.}\end{cases}$
@@ 3.1. $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
$f_{1}$ $=q+(\beta_{0})q^{2}+\cdots,$
$f_{2}$ $=q+(-\beta_{0}+1080)q^{2}+\cdots,$
@@
$f_{1}\equiv f_{2}\pmod{\mathfrak{P}_{144169}}$,
The space $S_{l}(\Gamma_{0}(N))=S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ is of dimension $1$ and generated by the unique Ramanujan cusp form $\Delta$. Put $g=\Delta.$ The Fourier expansion of $g$ is $q-24q^{2}+\cdots$.
@@ 3.1.5. The critical point $m=18$
$E_{2,1}^{*}(z,1\!\!1)=\frac{-12}{4\pi y}+1-24q-72q^{2}$-$\cdots$
See Miyake [9, Chapter 7]. By using Lemma 2.6 the holomorphic projection $h_{0}$ to be
@@ 3.2. $S_{30}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
where $\nu_{1}=\sqrt{-8424}$. It can checked that $f_{2}$ and $f_{3}$ are Galois conjugates of each other, and in fact, $f_{2\rho}=f_{2}^{\rho}=f_{3}$ and $f_{3\rho}=f_{3}^{\rho}=f_{2}$ as the coefficients are in the imaginary quadratic field $\mathbb{Q}(\nu_{1}).$ However, $f_{1}$ and $f_{2}$ are not Galois conjugate of each other, since the coefficients of $f_{1}$ are in $\mathbb{Q}$; also, $f_{1\rho}=f_{1}^{\rho}=f_{1},$ Using the Sturm bound (Thm. 2.10) one verifies
@@ 3.4. $S_{26}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{13}(\Gamma_{0}(3),\psi)$
If $h_{0}$ is the holomorphic projection and $h_{0}=c_{0}f+c_{1}\hat{f}+\cdots\in S_{26}(\Gamma_{1}(3))$ then it is no longer true that $\langle f,h_{0}\rangle=c_{0}\langle f,f\rangle,$ since $\langle f,\hat{f}\rangle\neq 0.$ Hence $c_{r}\alpha_{0}$ is not the correct $L$-value. So while returning the $L$-value this has to be taken into account. The relation between $\langle f,f\rangle$ and $\langle f,\hat{f}\rangle$ is as follows: Suppose $f\in S_{k}(\Gamma_{1}(N))$ be a newform. Let $a(p,f)$ be its coeffecient at $p$. In $S_{k}(\Gamma_{0}(Np)),$ taking $\hat{f}(z)=f(pz)$ one has:
@@ 3.5. $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))\times M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
We may also vary the modular forms in a congruence class with one being cuspidal and the other an Eisenstein series. In this section, we vary the lower weight modular forms of weight $l=12$, with one form being the Ramanujan $\Delta$-function, and the other form the weight $12$ Eisenstein series. More precisely, let $f$ be the newform of weight $k=24$ for $SL_{2}(\mathbb{Z})$ with $q$-expansion $f(q)=q+\beta_{0}q^{2}+\left(-48\beta_{0}+195660\right)q^{3}+O(q^{4}),$ where $\beta_{0}$ satisfies the polynomial $x^{2}-1080x-20468736\in\mathbb{Q}[x]$. This is one of the newforms from Section 3.1. For the smaller weight $l=12$, take $g_{1}:=\Delta\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g_{2}:=E_{12}(z)={691/65520}+q+\cdots\in M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$, the Eisenstein series of weight $12$ and level $1$. The well-known Ramanujan congruence asserts that $g_{1}\equiv g_{2}\pmod{691}$.
@@
In a similar vein, by Shimura [15, Lem. 1], one has:
@@
The first term inside $\left(\cdots\right)$ is rational due to well known result on the special values of the Riemann zeta function. The second and the third term are algebraic due to equations (11) and (12). Hence we use the algorithm described in Section 2.3.2 to calculate the ratios of the special values $\frac{L(m,f\times g_{2})}{L(m+1,f\times g_{2})}.$ The ratios $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ and $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ depending on the pairity of $m$ are summarized in the table.

### v2

@@ 2.3.2. Some preliminaries for $\mathrm{SL}_{2}(\mathbb{Z})$
    Find a basis $b_{1}^{-},b_{2}^{-},\dots,b_{d}^{-}$ for the space $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}=\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}.$
@@
    Express the Manin symbols $(i,0,1)$ for $i$ odd and $0\leq i\leq k-3$ in terms of the basis $b_{1}^{-},b_{2}^{-},\dots,b_{d}^{-}.$ Let $A^{-}=[a^{-}_{ij}]$ be the matrix with rational entries such that $(i,0,1)=\sum_{j=1}^{d}a_{ji}^{-}b_{j}^{-}.$
@@
    Compute the matrices $M_{2}^{-},\dots,M_{d}^{-}$ for the action of the Hecke operators $T_{2}^{-},\dots,T_{d}^{-}$ on $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}$ with respect to the basis $b_{1}^{-},b_{2}^{-},\dots,b_{d}^{-}.$
@@
    Form the matrix
    $\mathcal{L}^{-}=\left(\begin{array}[]{c}{{}^{t}}M_{2}^{-}-a(2,f)I_{d}\\ \vdots\\ {{}^{t}}M_{d}^{-}-a(d,f)I_{d}\end{array}\right).$
@@
[…]
@@
    Compute the $\frac{k-2}{2}\times 1$ column matrix $E:={{}^{t}}A^{-}\cdot w^{-}.$ Let $c_{1},c_{3},\dots,c_{k-3}$ be the entries of $E$.
  7. Step 7.
    $\textbf{Return:}\qquad\dfrac{1}{(i-1)!}\cdot\dfrac{c_{i}}{c_{1}}\text{ for $i=1,3,\dots,k-3$}.$
This is the desired output by (12) and
$\frac{(f,X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\})}{(f,Y^{k-2}\otimes\{0,\infty\})}=\frac{c_{i}}{c_{1}}\>\>\>\mbox{for}\>\>i=1,3,\dots k-3.$
See 2.3.5 for a proof of the algorithm.
2.3.4.
The algorithm for odd critical points
Input: The first $d$ Fourier coefficients $a(1,f)=1,\,a(2,f),\dots,\,a(d,f)$ of a normalized eigenform $f\in S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ which are assumed to be real.
Output: $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ for $m=1,3,\dots,k-1.$
  1. Step 1.
    Find a basis $b_{1}^{+},b_{2}^{+},\dots,b_{d+1}^{+}$ for the space $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{+}.$
  2. Step 2.
    Express the Manin symbols $(i,0,1)$ for $i$ even and $0\leq i\leq k-2$ in terms of the basis $b_{1}^{+},b_{2}^{+},\dots,b_{d+1}^{+}.$ Let $A^{+}=[a^{+}_{ij}$] be a matrix with rational entries such that $(i,0,1)=\sum_{j=1}^{d+1}a^{+}_{ji}b_{j}^{+}.$
  3. Step 3.
    Compute the matrices $M_{2}^{+},\dots,M_{d}^{+}$ of the Hecke operators $T_{2}^{+},\dots,T_{d}^{+}$ acting on $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{+}$ with respect to $b_{1}^{+},b_{2}^{+},\dots,b_{d+1}^{+}.$
  4. Step 4.
    Form the block matrix
    $\mathcal{L}^{+}=\left(\begin{array}[]{c}{{}^{t}}M_{2}^{+}-a(2,f)I_{d+1}\\ \vdots\\ {{}^{t}}M_{d}^{+}-a(d,f)I_{d+1}\end{array}\right).$
  5. Step 5.
[…]
  6. Step 6.
    Compute the $\frac{k}{2}\times 1$ column matrix $O:={{}^{t}}A^{+}\cdot w^{+}.$ Let $d_{0},d_{2},\dots,d_{k-2}$ be the entries of $O$.
  7. Step 7.
    $\textbf{Return:}\qquad\dfrac{1}{(i-1)!}\cdot\dfrac{d_{i}}{d_{0}}\text{ for $i=0,2,\dots,k-2$}.$
This is the desired output by (11) and
$\frac{(f,X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\})}{(f,Y^{k-3}\otimes\{0,\infty\})}=\frac{d_{i}}{d_{0}}\>\>\>\mbox{for}\>\>i=0,2,\dots,k-2.$
2.3.5. The correctness of the algorithms
The correctness of both the algorithms, after a slight modification, are the same as in Manin [8, Sections 4.1, 4.2, and 6]. For the reader’s convenience a proof of the algorithm is adumbrated for even critical points.
Let $\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ denote the $\mathbb{C}$-dual space of $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}.$ Define the Hecke action on the dual space $\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ in the usual way, i.e., the $n$-th Hecke operator $\tilde{T}^{-}_{n}$ acts on $\tilde{f}\in\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ by $\langle\!\langle f,\tilde{T}^{-}_{n}(\tilde{f})\rangle\!\rangle=\langle\!\langle T^{-}_{n}(f),\tilde{f}\rangle\!\rangle$, where $\langle\!\langle f,\tilde{f}\rangle\!\rangle$ is evaluation of $\tilde{f}$ on $f.$
Proof.
(For even critical points.) Since the $T_{n}$’s are a commuting family of diagonalizable operators, the space $S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ can be decomposed into Hecke isotypic subspaces which are all one-dimensional by multiplicity one. Since the pairing $(\cdot,\cdot)$ is Hecke equivariant, the same is true for the spaces $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ and $\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$. We can write
$\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}=\bigoplus_{f\in S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))}\mathbb{C}\omega_{f}$
for Hecke isotypic decomposition where $\omega_{f}(\_)=(f,\_)\in\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ has the same system of eigenvalues as $f$. The space $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ has a rational subspace $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}$ which is stable under the Hecke operators. Let $b_{1}^{-},b_{2}^{-},\dots,b_{d}^{-}$ be a basis of $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}$ and $\tilde{b}_{1}^{-},\tilde{b}_{2}^{-},\dots,\tilde{b}_{d}^{-}$ be the dual basis. Let $M_{n}^{-}$ be the $d\times d$ matrix of $T_{n}$ with respect to this basis. Then with respect to the dual basis $\tilde{b}_{1}^{-},\tilde{b}_{2}^{-},\dots,\tilde{b}_{d}^{-}$ the matrix of $\tilde{T}_{n}^{-}$ is simply ${{}^{t}}M_{n}^{-}$.
A vector in the isotypic subspace $\mathbb{C}\omega_{f}$ in $\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}$ lies in $\bigcap_{n}\text{Null}({{}^{t}}M_{n}^{-}-a(n,f)1\!\!1_{d}).$ Since $f$ is completely determined the first $d$ many Fourier coefficients, the intersection can be taken over a smaller set
$\bigcap_{n=1}^{d}\text{Null}({{}^{t}}M_{n}^{-}-a(n,f)1\!\!1_{d}).$
(This is Step 5 in the Algorithm where one row-reduces the matrix $\mathcal{L}^{-}$ to pick a vector from its null space.) Let $\tilde{w}^{-}={{}^{t}}(w_{1}^{-},w_{2}^{-},\dots,w_{n}^{-})=\sum_{i=1}^{d}w_{i}^{-}\tilde{b}_{i}^{-}$ be a non-zero vector from the intersection. By construction $\tilde{w}^{-}$ lies in the isotypic subspace of $\widetilde{\mathbb{S}}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{C})^{-}$ determined by $\tilde{T}_{n}^{-}\tilde{w}^{-}=a(n,f)\tilde{w}^{-}$. By multiplicity one, this is the same as $\mathbb{C}\omega_{f}$. Hence,
(13) $\tilde{w}^{-}=\lambda\omega_{f},\quad\mbox{for some $\lambda\in\mathbb{C}^{\times}$}.$
For an odd integer $i$ and $0\leq i\leq k-2,$ if $X^{i}Y^{k-2-i}\otimes\{0,\infty\}=\sum_{j=1}^{d}a_{ji}^{-}b_{j}^{-}$ with $a_{ij}\in\mathbb{Q}$ then
$\tilde{w}^{-}(X^{i}Y^{k-2-i}\otimes\{0,\infty\})=\tilde{w}^{-}(\sum_{j=1}^{d}a_{ji}^{-}b_{j}^{-})=\sum_{j=1}^{d}a_{ji}^{-}\tilde{w}^{-}(b_{j}^{-})=\sum_{j=1}^{d}a_{ji}^{-}w_{j}^{-}.$
Set $E:={{}^{t}}A^{-}\cdot w^{-}$, where $A^{-}=[a_{ij}^{-}]$, then from (13) it is clear that
${{}^{t}}E=\left(\lambda(f,X^{1}Y^{k-3}\otimes\{0,\infty\}),\,\,\lambda(f,X^{3}Y^{k-5}\otimes\{0,\infty\}),\dots,\,\,\lambda(f,X^{k-3}Y^{1}\otimes\{0,\infty\})\right)$
which proves the given algorithm. ∎
@@ 2.4. Sturm’s criterion for congruence between cusp forms
To check congruence between cusp forms we will use the following criterion of Sturm [17]; see also Stein [16, p. 173].
@@
Let $\mathfrak{P}$ be a prime ideal of a number field $K$. Let $f,g\in M_{k}(\Gamma,\mathcal{O}_{K})$ be modular forms of weight $k$ for a congruence subgroup $\Gamma$ of level $N$ with Fourier coefficients in the ring of integers $\mathcal{O}_{K}$ of $K$. Let $m=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma].$ Suppose
@@
$n\leq\begin{cases}\frac{km}{12}-\frac{m-1}{N}&\mbox{if $f-g\in S_{k}(\Gamma,\mathcal{O}_{K}$}),\\ \frac{km}{12}&\mbox{otherwise.}\end{cases}$
@@ 3.1. $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
$f_{1}$ $=q+(\beta_{0})q^{2}+O(q^{3}),$
$f_{2}$ $=q+(-\beta_{0}+1080)q^{2}+O(q^{3}),$
@@
$f_{1}\equiv f_{2}\pmod{\mathfrak{P}_{144169}}$.
The space $S_{l}(\Gamma_{0}(N))=S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ is of dimension $1$ and generated by the unique Ramanujan cusp form $\Delta$. Put $g=\Delta.$ The Fourier expansion of $g$ is $q-24q^{2}+O(q^{3})$.
@@ 3.1.5. The critical point $m=18$
$E_{2,1}^{*}(z,1\!\!1)=\frac{-12}{4\pi y}+1-24q-72q^{2}-\cdots.$
See Miyake [9, Chapter 7]. By using Lemma 2.6 the holomorphic projection $h_{0}$ is
@@ 3.2. $S_{30}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
where $\nu_{1}=\sqrt{-8424}$. It can checked that $f_{2}$ and $f_{3}$ are Galois conjugates of each other, and in fact, $f_{2\rho}=f_{2}^{\rho}=f_{3}$ and $f_{3\rho}=f_{3}^{\rho}=f_{2}$ as the coefficients are in the imaginary quadratic field $\mathbb{Q}(\nu_{1}).$ However, $f_{1}$ and $f_{2}$ are not Galois conjugate of each other, since the coefficients of $f_{1}$ are in $\mathbb{Q}$; also, $f_{1\rho}=f_{1}^{\rho}=f_{1},$ Using the Sturm bound (Thm. 2.11) one verifies
@@ 3.4. $S_{26}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{13}(\Gamma_{0}(3),\psi)$
If $h_{0}$ is the holomorphic projection and $h_{0}=\alpha_{0}f+\alpha_{1}\hat{f}+\dots\in S_{26}(\Gamma_{1}(3))$ then it is no longer true that $\langle f,h_{0}\rangle=c_{0}\langle f,f\rangle,$ since $\langle f,\hat{f}\rangle\neq 0.$ Hence $c_{r}\alpha_{0}$ is not the correct $L$-value. So while returning the $L$-value this has to be taken into account. The relation between $\langle f,f\rangle$ and $\langle f,\hat{f}\rangle$ is as follows: Suppose $f\in S_{k}(\Gamma_{1}(N))$ be a newform. Let $a(p,f)$ be its coeffecient at $p$. In $S_{k}(\Gamma_{0}(Np)),$ taking $\hat{f}(z)=f(pz)$ one has:
@@ 3.5. $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))\times M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
We may also vary the modular forms in a congruence class with one being cuspidal and the other an Eisenstein series. In this section, we vary the lower weight modular forms of weight $l=12$, with one form being the Ramanujan $\Delta$-function, and the other form the weight $12$ Eisenstein series. More precisely, let $f$ be the newform of weight $k=24$ for $\mathrm{SL}_{2}(\mathbb{Z})$ with $q$-expansion $f(q)=q+\beta_{0}q^{2}+\left(-48\beta_{0}+195660\right)q^{3}+O(q^{4}),$ where $\beta_{0}$ satisfies the polynomial $x^{2}-1080x-20468736\in\mathbb{Q}[x]$. This is one of the newforms from Section 3.1. For the smaller weight $l=12$, take $g_{1}:=\Delta\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g_{2}:=E_{12}(z)={691/65520}+q+\cdots\in M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$, the Eisenstein series of weight $12$ and level $1$. The well-known Ramanujan congruence asserts that $g_{1}\equiv g_{2}\pmod{691}$.
@@
In a similar vein, by Shimura [14, Lem. 1], one has:
@@
The first term inside $\left(\cdots\right)$ is rational due to well known result on the special values of the Riemann zeta function. The second and the third term are algebraic due to equations (11) and (12). Also, the reader is reminded that the form $f$ has totally real coefficients. Hence we use the algorithms described in 2.3.3 and 2.3.4 to calculate the ratios of the special values $\frac{L(m,f\times g_{2})}{L(m+1,f\times g_{2})}.$ The ratios $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ and $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ depending on the pairity of $m$ are summarized in the table.